Upon graduating with an accounting degree, you open your own accounting firm of which you are the sole employee. To start the firm you passed on a job offer with a large accounting firm that offered you a salary of RM60,000 annually. Last year you earned a total revenue of RM100,000. Rent and supplies last year were RM50,000.

 

2.1) Your annual economic costs are RM110,000 . (1M)

 

2.2) Your annual economic profit is RM-10,000. (1M)

 

2.3) Your annual accounting profit is RM50,000 (1M)

 

2.4) Based on the above information, describe the condition of your firm. (2M)

_________________________________________________________________________________________________________________________________________________________________________________________________________________________________


1
Expert's answer
2021-06-21T16:11:37-0400

Solution

a.) EC= 50,000+60,000=110,000


b.)EP= TR-TEC

100,000-110,000= -10,000


c.)AP=TR-TC

100,000-50,000=50,000


d.) Based on the above information,the firm is operating on losses since there's no profit gained.